# CERN-TH-2023-230 WUB/23-02 HU-EP-23/72 Heavy Wilson Quarks and O($a$) Improvement: Nonperturbative Results for $b_{\rm g}$

### 30 Dec 2023 v1

@@ Abstract
With Wilson quarks, on-shell O($a$) improvement of the lattice QCD action is achieved by including the Sheikholeslami-Wohlert term and two further operators of mass dimension 5, which amount to a mass-dependent rescaling of the bare parameters. We here focus on the rescaled bare coupling, ${\tilde{g}_{0}^{2}}=g_{0}^{2}(1+b_{\rm g}am_{\rm q})$, and the determination of $b_{\rm g}(g_{0}^{2})$, which is currently only known to 1-loop order of perturbation theory. We derive suitable improvement conditions in the chiral limit and in a finite space-time volume and evaluate these for different pure gauge observables, both with and without the gradient flow. The choice of $\beta$-values and the line of constant physics are motivated by the ALPHA collaboration’s decoupling strategy to determine $\alpha_{s}(m_{Z})$ [1]. However, the improvement conditions and some insight into systematic effects may prove useful in other contexts, too.
@@ 1 Introduction
Various $b$-coefficients have been determined non-perturbatively over the past 25 years, including $b_{\mathrm{m}}$ and some of the coefficients required for the O($a$) improvement of quark bilinear composite operators (see [6, 7, 8, 9, 10, 11, 12, 13] for an incomplete list of references). The coefficient $b_{\rm g}$ is a notable exception. While there exist some qualitative ideas in the literature as to how $b_{\rm g}$ could be determined by measurement (see e.g. [14]), no proof of concept regarding their viability and practicality has been given. This is somewhat surprising given the central rôle of $b_{\rm g}$ for the consistency of O($a$) improvement: it is required when the quark masses are varied at fixed lattice spacing such as needed for chiral extrapolations or in studies of decoupling [5, 1]. Furthermore, since constant lattice spacing amounts to keeping ${\tilde{g}_{0}^{2}}$ constant, the renormalization constants in mass-independent schemes depend on ${\tilde{g}_{0}^{2}}$ rather than $g_{0}^{2}$. In fact, for composite operators this often leads to the determination of effective $b$-coefficients, which include the $b_{\rm g}$-contribution from the renormalization constant by a Taylor expansion to first order in $am_{\rm q}$ [11], even if not mentioned explicitly [12].
@@ 2.1 Chiral Ward identities and $b_{\rm g}$
and $d_{\rm S}$ can, in prinicple, be determined by chiral Ward identities, together with $c_{\rm S}$ and the scale independent ratio $Z_{\rm P}/Z_{\rm S^{0}}$ [20]. A difficulty is posed by the fact that the chiral limit is only defined up to cutoff effects of O($a^{2}$), which render the subtraction of the cubic power divergence ambiguous. More precisely, with $c_{\rm S}\equiv c_{\rm S}(g_{0}^{2},am_{\rm q})$, an expansion in powers of $am_{\rm q}$ yields a term $\propto am_{\rm q}/a^{3}=m_{\rm q}/a^{2}$. Given the intrinisic O($a^{2}$) ambiguity of $m_{\rm q}$, this results in an ambiguity of O($1$) in the subtraction term rendering the definition of the renormalized scalar density impossible along these lines. It is therefore important to note that one may combine Ward identities such that only connected correlation functions remain, in which all power divergences cancel. A practical implementation using SF boundary conditions [21, 22] can be obtained following the discussion in [20]. Whether these observations lead to a practical method for a non-perturbative determination of $b_{\rm g}$ remains to be seen. We emphasize that this strategy relies entirely on the identity (2.4). We would therefore like to point out a subtlety that was overlooked in [19]. Given that one relates an improved operator to the improved action parameters, it is natural to expect that the O($a$) counterterm in Eq. (2.3) cannot be discretized arbitrarily, but must be related to the derivative of the lattice action density with respect to $g_{0}^{2}$. For definiteness, the O($a$) improved lattice action
@@ 2.2 Restoration of chiral symmetry in a small volume
In this paper we pursue a different approach. It is based on the observation that, in the absence of spontaneous chiral symmetry breaking, physical quark mass effects in gluonic observables are quadratic or higher order in the quark mass. In contrast, the counterterm proportional to $b_{\rm g}$ is designed to cancel terms that are linear in the quark mass, allowing us, in principle, to distinguish these effects. Let us first have a closer look at the physical quark mass dependence. If we consider a finite space time manifold without boundaries, the lattice QCD partition function with Ginsparg-Wilson type quarks becomes a finite dimensional, mathematically well-defined ordinary and Grassmann-integral, with exact chiral and flavour symmetries. The absence of spontaneous symmetry breaking in a finite volume, implies that the partition function is an analytic function of $m_{\rm q}$. For even $N_{\mathrm{f}}$, a change of the fermionic variables by a discrete chiral field transformation [^2] On the lattice with Ginsparg-Wilson quarks and Neuberger’s ($\gamma_{5}$-hermitian) Dirac operator $D_{N}$ [23], the first of the $\gamma_{5}$ factors is replaced by $\hat{\gamma}_{5}\equiv\gamma_{5}(1-aD_{N})$ [24]. Note that, for even $N_{\mathrm{f}}$, the transformation is part of the non-singlet chiral symmetry.
@@
then establishes that this function must be even, implying the absence of terms linear in $m_{\rm q}$ for any gluonic observable. With odd $N_{\mathrm{f}}>2$, this is no longer true; however, by adapting the argument of appendix D in ref. [25] to the lattice regularization with Ginsparg-Wilson quarks [26, 27, 23, 28, 24], the discrete chiral field transformation,
@@
leads to a change of variables with unit Jacobian that leaves the massless action invariant. Under the same transformation, a single insertion of the flavour singlet scalar density, $S^{0}$, into a gluonic correlation function is proportional to itself and must vanish, provided that both the QCD action and the gluonic observable are parity even [^3] We again use a continuum notation; on the lattice, besides using $\hat{\gamma}_{5}$ [24] for the transformation of $\psi$, the scalar lattice density takes the form $S^{0}=\overline{\psi}(1-\frac{a}{2}D_{N})\psi$, such that it correctly transforms under the lattice chiral symmetry (see e.g. Section 2 in [29]).. Although this does not exclude all odd powers in the quark mass, it means that contributions linear in the quark mass are indeed absent.
We will denote a generic renormalized gluonic observable by $O_{g}$. Gradient flow observables, such as the gauge action density at finite flow time [30], are natural candidates, but the gradient flow is not a necessary requirement. We choose a finite Euclidean space time volume $L^{4}$ with boundary conditions that do not break all the chiral symmetries. A hyper-torus geometry with some kind of periodic boundary conditions for all fields is an obvious possibility, however, other possible options include chirally rotated SF ($\chi$SF) boundary conditions (with even $N_{\mathrm{f}}$[31, 32], or a mixture of SF [21, 22] and $\chi$SF boundary conditions (with odd $N_{\mathrm{f}}$[33]. For simplicity we also assume that $O_{g}$ has no mass dimension, which can always be achieved by multiplication with the appropriate power of $L$. In the continuum limit, the observable thus becomes a function of two dimensionless variables, which we may choose as $z=ML$ and $\Lambda_{\rm\overline{MS\kern-0.35002pt}\kern 0.35002pt}^{(3)}L$, with $M$ and $\Lambda_{\rm\overline{MS\kern-0.35002pt}\kern 0.35002pt}^{(3)}\equiv\Lambda$ being the Renormalization Group Invariant (RGI) quark mass and $\Lambda$-parameter of the $N_{\mathrm{f}}=3$ theory, respectively. From the previous discussion we then expect that the continuum limit of the lattice expectation value, $\langle O_{g}\rangle$, becomes a function of $(z,\Lambda L)$, with an expansion in the RGI mass of the form
@@ 3.2 Creutz ratios
where the lattice derivatives ~$\tilde{\partial}_{y}f(y)=[f(y+a)-f(y-a)]/(2a)$ are symmetric to avoid linear terms in $a$. These Creutz ratios are made dimensionless and restricted to the diagonal ones,
@@
We have fixed the ratio $R/L$ to $1/4$ since this is a practical number for lattices with $L/a=12,16,\ldots$, which we will use below. Also $\hat{\chi}$ is a finite pure gauge observable for the following reasons. In continuous space-time, Wilson loops around smooth paths have been shown to be finite up to an overall factor $Z_{W}(l)$ which depends (apart from coupling and regularisation parameter) only on the the length $l$ of the path [40]. When the path contains cusps, such as the rectangular Wilson loops, there is an additional renormalization factor depending on the angles of the cusps. Both factors are removed by the derivatives in eq. 3.15. In our implementation on the lattice, we apply one HYP2 smearing step with parameters $\alpha_{1}=1,\,\alpha_{2}=1$ and $\alpha_{3}=0.5$ to the link variables which form the loop [41, 42, 43, 44, 45]. Since this does not change the symmetry properties of the links, $\hat{\chi}$ remains finite and has the same continuum limit as the observable with no smearing. Finally we need also the absence of terms linear in $a$ in the observable. Their absence for the static quark potential has been explained in [46]. Analogously it holds for Creutz ratios defined with symmetric derivatives.
@@ 4.2 Simulations
The simulation setup is similar to previous simulations with $N_{\mathrm{f}}=3$ degenerate quarks. We use an even-odd preconditioned variant of the HMC algorithm [47] in which the three quarks are separated into a doublet and a third quark. The quark determinant for the doublet is factorized [48] into three factors $\det[\hat{D}^{\dagger}\hat{D}]=\det[\hat{D}^{\dagger}\hat{D}+\mu_{2}^{2}]\times\det[(\hat{D}^{\dagger}\hat{D}+\mu_{1}^{2})(\hat{D}^{\dagger}\hat{D}+\mu_{2}^{2})^{-1}]\times\det[(\hat{D}^{\dagger}\hat{D})(\hat{D}^{\dagger}\hat{D}+\mu_{1}^{2})^{-1}]$ and pseudo fermion fields are introduced for each factor. Here $\hat{D}$ denotes the even-odd preconditioned clover Wilson operator. For the third quark it is required that all eigenvalues of $\hat{D}$ have real parts that are larger than some $r_{a}>0$. The operator in the determinant $\det[\hat{D}]=\det[\sqrt{\hat{D}^{\dagger}\hat{D}}]$ can then be approximated by a rational function of $\hat{D}^{\dagger}\hat{D}$. The inexactness of this approximation is accounted for by including a stochastically determined reweighting factor in the observables. The molecular dynamics equations derived from the various contributions to the action are solved using multi level integration schemes with two levels. The gauge action is integrated with the finest resolution while forces from different pseudo-fermion actions are evaluated less frequently. We use a modified version of openQCD-1.6 [49] for our simulations and refer to its documentation for further details. Simulation parameters as well as tests concerning the spectral gap and the ergodicity of the simulations are collected in appendix B.
@@ 5 Data analysis
The data for $\hat{\chi}$ are analyzed in the same way. Results for $b_{\rm g}$ from both observables are listed in table 3. As checks on the dependence on $c$ and $L/a$-values deviating from the lines of constant physics, there are additional data not reported in the tables.
@@
Figure 3: $b_{\rm g}$, eq. 2.13, from $\langle O_{g}\rangle=R^{2}\chi(R,R)$ vs. $R/a$ at $\beta=4.9$ for $L/a=16$, $24$, and $32$ (left) and $\beta=4.3020$, $4.4662$ and $4.7141$ for $L/a=12$, $16$, $24$ (right).
@@ 7 Conclusions
We emphasize that eq. 6.20 has not been shown to be valid below $\beta=4.3$. Firstly, we simulated only starting at that $\beta$-value, and secondly the LCP would quickly demand very small $\sqrt{8t}/a$ where ambiguities are potentially very large. Extending the present method to the region of CLS simulations [17, 18] is therefore not entirely straightforward. One alternative are higher order perturbative estimates; for recent steps towards a 2-loop result cf. [50], and one of us (MDB) has derived a preliminary result for the Wilson gauge action using the code developped for the 2-loop calculation of the SF coupling [51, 52]. Another alternative is the chiral Ward identity technique, applied to the O($a$) improved flavour singlet scalar density, eq. (2.3), cf. [20]. The connection to $b_{\rm g}$ relies on the identity, eq. (2.4), which was first derived by Bhattacharya et al. in [19]. In Section 2, we have pointed out an oversight in this original derivation. The identity can however be rescued, provided one uses the particular discretization of the O($a$) counterterm to the scalar density, eq. (2.7), which includes a contribution from the Sheikholeslami-Wohlert term.
@@ Appendix B Simulations
The factorization of the determinant of the doublet into three determinants, as described in section 4.2, has in all cases the masses $a\mu_{1}=0.1$ and $a\mu_{2}=1.5$. The degrees and ranges of the rational approximations for the third quark’s determinant vary and can be found in table 4. This table also shows the acceptance rates. For solving linear systems, either a standard conjugate gradient solver, or a variant that solves systems with different shifts simultaneously, is used and stopped when a relative residuum norm below $10^{-12}$ is reached.
@@
Figure 5: The lowest eigenvalue of $|\gamma_{5}\hat{D}|$ on several thermalized configurations at two different lattice spacings. The upper part corresponds to the finer spacing ($L/a=24,\beta=4.7141$), the lower one to the coarser ($L/a=16,\beta=4.4662$). Both are at approximately zero quark mass. The vertical red lines correspond to the lower interval boundaries of the rational approximations that were used in the simulations.
Data analysis is performed using ADerrors.jl [54, 55] for the gradient flow and pyerrors [56] for the Wilson loops keeping track of the correct errors and auto-correlations via the $\Gamma$-method [57, 58].
At last a remark on the ergodicity of the HMC algorithm. There are two potential difficulties with our setup. First, it is known [58] that with decreasing lattice spacing the correct sampling of topological sectors becomes increasingly difficult. This does not affect our calculation because at the volume corresponding to our LCP the sector with non-zero charge are suppressed so strongly, that they play virtually no rôle. The second is a difficulty related to the center symmetry. In pure gauge theory, in the large volume limit, the center symmetry is spontaneously broken at high enough temperature. In our small volume the symmetry is restored but the trace of a Polyakov line exhibits a tri-modal distribution and many algorithms, including the HMC, tend to get stuck in one of the sectors. Although there is no center symmetry to brake in the case of full QCD, similar algorithmic issues have been observed. We investigated the distribution of Polyakov lines and observed that at the volume corresponding to our LCP the distribution is unimodal, and in the perturbative region (e.g. $L/a=12,\beta=16$) there is enough movement between the “sectors” to ensure ergodicity.

### 01 Feb 2024 v2

# CERN-TH-2023-230 WUB/23-02 HU-EP-23/72 DESY-24-001 Heavy Wilson Quarks and O($a$) Improvement: Nonperturbative Results for $b_{\rm g}$

@@ Abstract
With Wilson quarks, on-shell O($a$) improvement of the lattice QCD action is achieved by including the Sheikholeslami-Wohlert term and two further operators of mass dimension 5, which amount to a mass-dependent rescaling of the bare parameters. We here focus on the rescaled bare coupling, ${\tilde{g}_{0}^{2}}=g_{0}^{2}(1+b_{\rm g}am_{\rm q})$, and the determination of $b_{\rm g}(g_{0}^{2})$, which is currently only known to 1-loop order of perturbation theory. We derive suitable improvement conditions in the chiral limit and in a finite space-time volume and evaluate these for different gluonic observables, both with and without the gradient flow. The choice of $\beta$-values and the line of constant physics are motivated by the ALPHA collaboration’s decoupling strategy to determine $\alpha_{s}(m_{Z})$ [1]. However, the improvement conditions and some insight into systematic effects may prove useful in other contexts, too.
@@ 1 Introduction
Various $b$-coefficients have been determined non-perturbatively over the past 25 years, including $b_{\mathrm{m}}$ and some of the coefficients required for the O($a$) improvement of quark bilinear composite operators (see [6, 7, 8, 9, 10, 11, 12, 13, 14] for an incomplete list of references). The coefficient $b_{\rm g}$ is a notable exception. While there exist some qualitative ideas in the literature as to how $b_{\rm g}$ could be determined by measurement (see e.g. [15]), no proof of concept regarding their viability and practicality has been given. This is somewhat surprising given the central rôle of $b_{\rm g}$ for the consistency of O($a$) improvement: it is required when the quark masses are varied at fixed lattice spacing such as needed for chiral extrapolations or in studies of decoupling [5, 1]. Furthermore, since constant lattice spacing amounts to keeping ${\tilde{g}_{0}^{2}}$ constant, the renormalization constants in mass-independent schemes depend on ${\tilde{g}_{0}^{2}}$ rather than $g_{0}^{2}$. In fact, for composite operators this often leads to the determination of effective $b$-coefficients, which include the $b_{\rm g}$-contribution from the renormalization constant by a Taylor expansion to first order in $am_{\rm q}$ [11], even if not mentioned explicitly [12].
@@ 2.1 Chiral Ward identities and $b_{\rm g}$
and $d_{\rm S}$ can, in prinicple, be determined by chiral Ward identities, together with $c_{\rm S}$ and the scale independent ratio $Z_{\rm P}/Z_{\rm S^{0}}$ [21]. A difficulty is posed by the fact that the chiral limit is only defined up to cutoff effects of O($a^{2}$), which render the subtraction of the cubic power divergence ambiguous. More precisely, with $c_{\rm S}\equiv c_{\rm S}(g_{0}^{2},am_{\rm q})$, an expansion in powers of $am_{\rm q}$ yields a term $\propto am_{\rm q}/a^{3}=m_{\rm q}/a^{2}$. Given the intrinsic O($a^{2}$) ambiguity of $m_{\rm q}$, this results in an ambiguity of O($1$) in the subtraction term rendering the definition of the renormalized scalar density impossible along these lines. It is therefore important to note that one may combine Ward identities such that only connected correlation functions remain, in which all power divergences cancel. A practical implementation using SF boundary conditions [22, 23] can be obtained following the discussion in [21]. Whether these observations lead to a practical method for a non-perturbative determination of $b_{\rm g}$ remains to be seen. We emphasize that this strategy relies entirely on the identity (2.4). We would therefore like to point out a subtlety that was overlooked in [20]. Given that one relates an improved operator to the improved action parameters, it is natural to expect that the O($a$) counterterm in Eq. (2.3) cannot be discretized arbitrarily, but must be related to the derivative of the lattice action density with respect to $g_{0}^{2}$. For definiteness, the O($a$) improved lattice action
@@ 2.2 Restoration of chiral symmetry in a small volume
In this paper we pursue a different approach. It is based on the observation that, in the absence of spontaneous chiral symmetry breaking, physical quark mass effects in gluonic observables are quadratic or higher order in the quark mass. In contrast, the counterterm proportional to $b_{\rm g}$ is designed to cancel terms that are linear in the quark mass, allowing us, in principle, to distinguish these effects. Let us first have a closer look at the physical quark mass dependence. If we consider a finite space time manifold without boundaries, the lattice QCD partition function with Ginsparg-Wilson type quarks [24, 25, 26, 27, 28] becomes a finite dimensional, mathematically well-defined ordinary and Grassmann-integral, with exact chiral and flavour symmetries. The absence of spontaneous symmetry breaking in a finite volume, implies that the partition function is an analytic function of $m_{\rm q}$. For even $N_{\mathrm{f}}$, a change of the fermionic variables by a discrete chiral field transformation [^2] On the lattice with Ginsparg-Wilson quarks and Neuberger’s ($\gamma_{5}$-hermitian) Dirac operator $D_{N}$ [26], the first of the $\gamma_{5}$ factors is replaced by $\hat{\gamma}_{5}\equiv\gamma_{5}(1-aD_{N})$ [28]. Note that, for even $N_{\mathrm{f}}$, the transformation is part of the non-singlet chiral symmetry.
@@
then establishes that this function must be even, implying the absence of terms linear in $m_{\rm q}$ for any gluonic observable. With odd $N_{\mathrm{f}}>2$, this is no longer true; however, by adapting the argument of appendix D in ref. [29] to the lattice regularization with Ginsparg-Wilson quarks, the discrete chiral field transformation,
@@
leads to a change of variables with unit Jacobian that leaves the massless action invariant. Under the same transformation, a single insertion of the flavour singlet scalar density, $S^{0}$, into a gluonic correlation function is proportional to itself and must vanish, provided that both the QCD action and the gluonic observable are parity even. [^3] We again use a continuum notation; on the lattice, besides using $\hat{\gamma}_{5}$ [28] for the transformation of $\psi$, the scalar lattice density takes the form $S^{0}=\overline{\psi}(1-\frac{a}{2}D_{N})\psi$, such that it correctly transforms under the lattice chiral symmetry (see e.g. Section 2 in [30]). Although this does not exclude all odd powers in the quark mass, it means that contributions linear in the quark mass are indeed absent.
We will denote a generic renormalized gluonic observable by $O_{g}$. Gradient flow observables [31, 32], such as the gauge action density at finite flow time [32], are natural candidates, but the gradient flow is not a necessary requirement. We choose a finite Euclidean space time volume $L^{4}$ with boundary conditions that do not break all the chiral symmetries. A hyper-torus geometry with some kind of periodic boundary conditions for all fields is an obvious possibility, however, other possible options include chirally rotated SF ($\chi$SF) boundary conditions (with even $N_{\mathrm{f}}$[33, 34], or a mixture of SF [22, 23] and $\chi$SF boundary conditions (with odd $N_{\mathrm{f}}$[35]. For simplicity we also assume that $O_{g}$ has no mass dimension, which can always be achieved by multiplication with the appropriate power of $L$. In the continuum limit, the observable thus becomes a function of two dimensionless variables, which we may choose as $z=ML$ and $\Lambda_{\rm\overline{MS\kern-0.35002pt}\kern 0.35002pt}^{(3)}L$, with $M$ and $\Lambda_{\rm\overline{MS\kern-0.35002pt}\kern 0.35002pt}^{(3)}\equiv\Lambda$ being the Renormalization Group Invariant (RGI) quark mass and $\Lambda$-parameter of the $N_{\mathrm{f}}=3$ theory, respectively. From the previous discussion we then expect that the continuum limit of the lattice expectation value, $\langle O_{g}\rangle$, becomes a function of $(z,\Lambda L)$, with an expansion in the RGI mass of the form
@@ 3.2 Creutz ratios
where the lattice derivatives ~$\tilde{\partial}_{y}f(y)=[f(y+a)-f(y-a)]/(2a)$ are symmetric to avoid linear terms in $a$. Note that, analogously to the energy density at positive flow time in eq. (3.14), also the Wilson loops $W(R,T)$ defining the Creutz ratios are projected to the topologically trivial sector. These Creutz ratios are made dimensionless and restricted to the diagonal ones,
@@
We have fixed the ratio $R/L$ to $1/4$ since this is a practical number for lattices with $L/a=12,16,\ldots$, which we will use below. Also $\hat{\chi}$ is a finite pure gauge observable for the following reasons. In continuous space-time, Wilson loops around smooth paths have been shown to be finite up to an overall factor $Z_{W}(l)$ which depends (apart from coupling and regularisation parameter) only on the length $l$ of the path [42]. When the path contains cusps, such as the rectangular Wilson loops, there is an additional renormalization factor depending on the angles of the cusps. Both factors are removed by the derivatives in eq. 3.15. In our implementation on the lattice, we apply one HYP2 smearing step with parameters $\alpha_{1}=1,\,\alpha_{2}=1$ and $\alpha_{3}=0.5$ to the link variables which form the loop [43, 44, 45, 46, 47]. Since this does not change the symmetry properties of the links, $\hat{\chi}$ remains finite and has the same continuum limit as the observable with no smearing. Finally we need also the absence of terms linear in $a$ in the observable. Their absence for the static quark potential has been explained in [48]. Analogously it holds for Creutz ratios defined with symmetric derivatives.
@@ 4.2 Simulations
The simulation setup is similar to previous simulations with $N_{\mathrm{f}}=3$ degenerate quarks. We use an even-odd preconditioned variant of the HMC algorithm [49] in which the three quarks are separated into a doublet and a third quark. The quark determinant for the doublet is factorized [50] into three factors $\det[\hat{D}^{\dagger}\hat{D}]=\det[\hat{D}^{\dagger}\hat{D}+\mu_{2}^{2}]\times\det[(\hat{D}^{\dagger}\hat{D}+\mu_{1}^{2})(\hat{D}^{\dagger}\hat{D}+\mu_{2}^{2})^{-1}]\times\det[(\hat{D}^{\dagger}\hat{D})(\hat{D}^{\dagger}\hat{D}+\mu_{1}^{2})^{-1}]$ and pseudo fermion fields are introduced for each factor. Here $\hat{D}$ denotes the even-odd preconditioned clover Wilson operator. For the third quark it is required that all eigenvalues of $\hat{D}$ have real parts that are larger than some $r_{a}>0$. The operator in the determinant $\det[\hat{D}]=\det[\sqrt{\hat{D}^{\dagger}\hat{D}}]$ can then be approximated by a rational function of $\hat{D}^{\dagger}\hat{D}$ [51, 52]. The inexactness of this approximation is accounted for by including a stochastically determined reweighting factor in the observables. The molecular dynamics equations derived from the various contributions to the action are solved using multi level integration schemes with two levels [53]. The gauge action is integrated with the finest resolution while forces from different pseudo-fermion actions are evaluated less frequently. We use a modified version of openQCD-1.6 [54] for our simulations and refer to its documentation for further details. Simulation parameters as well as tests concerning the spectral gap and the ergodicity of the simulations are collected in appendix B.
@@ 5 Data analysis
The data for $\hat{\chi}$ are analyzed in the same way. Results for $b_{\rm g}$ from both observables are listed in table 3. As checks on the dependence on $c$, $R/L$, and $L/a$-values deviating from the lines of constant physics, there are additional data not reported in the tables.
@@
Figure 3: $b_{\rm g}$, eq. 2.13, from $\langle O_{g}\rangle=R^{2}\chi(R,R)$ vs. $R/a$ at $\beta=4.9$ for $L/a=16$, $24$, and $32$ (left) and $\beta=4.3020$, $4.4662$, $4.7141$ for $L/a=12$, $16$, and $24$ (right).
@@ 7 Conclusions
We emphasize that eq. 6.20 has not been shown to be valid below $\beta=4.3$. Firstly, we simulated only starting at that $\beta$-value, and secondly the LCP would quickly demand very small $\sqrt{8t}/a$ where ambiguities are potentially very large. Extending the present method to the region of CLS simulations [18, 19] is therefore not entirely straightforward. One alternative are higher order perturbative estimates; for recent steps towards a 2-loop result cf. [55], and one of us (MDB) has derived a preliminary result for the Wilson gauge action using the code developed for the 2-loop calculation of the SF coupling [56, 57]. Another alternative is the chiral Ward identity technique, applied to the O($a$) improved flavour singlet scalar density, eq. (2.3), cf. [21]. The connection to $b_{\rm g}$ relies on the identity, eq. (2.4), which was first derived by Bhattacharya et al. in [20]. In Section 2, we have pointed out an oversight in this original derivation. The identity can however be rescued, provided one uses the particular discretization of the O($a$) counterterm to the scalar density, eq. (2.7), which includes a contribution from the Sheikholeslami-Wohlert term.
@@ Appendix B Simulations
The factorization of the determinant of the doublet into three determinants, as described in section 4.2, has in all cases the masses $a\mu_{1}=0.1$ and $a\mu_{2}=1.5$. The degrees and ranges of the rational approximations for the third quark’s determinant vary and can be found in table 4. This table also shows the acceptance rates. For solving linear systems, either a standard conjugate gradient solver, or a variant that solves systems with different mass shifts simultaneously, is used and stopped when a relative residuum norm below $10^{-12}$ is reached.
@@
Figure 5: The lowest eigenvalue of $|\gamma_{5}\hat{D}|$ on several thermalized configurations at two different lattice spacings. The upper part corresponds to the finer lattice spacing ($L/a=24,\beta=4.7141$), while the lower one to the coarser one ($L/a=16,\beta=4.4662$). Both are at approximately zero quark mass. The vertical red lines correspond to the lower interval boundaries of the rational approximations that were used in the simulations.
Data analysis is performed using ADerrors.jl [59, 60] for the gradient flow energy density and pyerrors [61] for the Wilson loops, keeping track of the correct errors and auto-correlations via the $\Gamma$-method [62, 63].
At last a remark on the ergodicity of the HMC algorithm. There are two potential difficulties with our setup. First, it is known [63] that with decreasing lattice spacing the correct sampling of topological sectors becomes increasingly difficult. This does not affect our calculation because at the volume corresponding to our LCP the sectors with non-zero charge are suppressed so strongly, that they play virtually no rôle. The second is a difficulty related to the center symmetry. In pure gauge theory, in the large volume limit, the center symmetry is spontaneously broken at high enough temperature. In our small volume the symmetry is restored, but the trace of a Polyakov line exhibits a tri-modal distribution and many algorithms, including the HMC, tend to get stuck in one of the sectors. Although there is no center symmetry to break in the case of full QCD, similar algorithmic issues have been observed. We investigated the distribution of Polyakov lines and observed that at the volume corresponding to our LCP the distribution is unimodal, and in the perturbative region (e.g. $L/a=12,\beta=16$) there is enough movement between the “sectors” to ensure ergodicity.